# Pauli Propagation for Imaginary Time Evolution

### v1

@@ III Numerical experiments
In this section, we present numerical results illustrating the performance of ITPP for approximating ground-state properties of the one-dimensional transverse-field Ising model (TFIM) with open boundary conditions. The Hamiltonian is
$\hat{H}_{\mathrm{TFIM}}=-J\sum_{i=1}^{n-1}Z_{i}Z_{i+1}-h\sum_{i=1}^{n}X_{i},$ (18)
@@
All simulations were performed using the PauliPropagation.jl Julia package [rudolph2025pauli], which provides an efficient implementation of Pauli propagation for real-time dynamics. We extend this framework by incorporating the imaginary time propagation rules introduced in Sec. II.3.
Since the ordered phase ($J>h$) is known to be the more challenging regime for approximate classical methods, our numerical study focuses on the parameter choice $J=1.0$ and $h=0.5$. In this setting, we propagate the identity operator in imaginary time—following the prescription of Sec. II.3—and evaluate the converged ground-state energy under various truncation schemes.
We first verify that ITPP faithfully reproduces Trotterized imaginary time evolution and compare its performance against exact imaginary time evolution obtained via exact diagonalization. Simulations are carried out for a chain of $N=10$ spins and a Trotter step size $\Delta t=0.04$. Throughout the imaginary time evolution, we track the expectation value of the energy and the relative error with respect to the exact ground-state energy. The corresponding results are presented in Fig. 1.
From the results, we observe that ITPP exactly matches the Trotterized imaginary time evolution, as expected, and closely follows the exact imaginary time evolution until Trotter errors become significant.
We next assess the performance of ITPP under a truncation strategy based on the magnitude of Pauli coefficients. The algorithm is executed using truncation thresholds ranging from $2^{-14}$ to $2^{-5}$, along with the idealized case in which no truncation is applied. Simulations are carried out in the ordered phase of the transverse-field Ising model with parameters $J=1.0$ and $h=0.5$, for a chain of $N=12$ spins and a Trotter step size $\Delta t=0.04$. Throughout the imaginary time evolution, we track the instantaneous energy, the relative error with respect to the exact ground-state energy, and the number of Pauli terms retained by the algorithm. The corresponding results are presented in Fig. 2.
ITPP with fixed-$K$ truncation and increasing NN
Figure 3: Scalability of imaginary time Pauli Propagation with fixed-$K$ truncation in the ordered phase of the one-dimensional TFIM ($J=1$, $h=0.5$). The left panel shows the converged ground-state energy obtained with ITPP as a function of system size NN, while the right panel displays the corresponding relative error (on a logarithmic scale) with respect to the exact ground-state energy computed using the Bogoliubov–de Gennes solution. The truncation parameter is fixed to K=2,704,156K=2{,}704{,}156, equal to the number of Pauli terms generated for $N=12$ in the absence of truncation. The results indicate the range of system sizes for which this fixed Pauli basis size is sufficient to accurately capture the underlying operator growth under imaginary time evolution, as well as the onset of a breakdown when the fixed computational budget becomes inadequate.
@@
For thresholds larger than $2^{-7}$, however, we observe a drastic reduction in the number of Pauli terms retained during the evolution, which results in poor approximations of the true ground state. This behavior can be attributed to the tree-like structure underlying Pauli Propagation: during imaginary time evolution, Pauli strings can generate many descendants at later times. If certain branches of this propagation tree are truncated too early, the algorithm discards Pauli terms that would otherwise contribute significantly at later stages. As a result, the evolution becomes confined to an overly restricted operator basis, leading to a loss of accuracy in the final ground-state approximation.
Truncation based on coefficient magnitude provides a natural way to discard Pauli terms by retaining only those with sufficiently large weights. However, this strategy does not offer direct control over the computational cost: fixing a threshold a priori does not determine how many Pauli terms will be retained during the evolution. In practice, the number of surviving terms can vary significantly depending on the system size and the stage of imaginary time evolution.
@@
We apply this fixed-$K$ truncation after each imaginary time operator. Since each operation can generate at most twice as many Pauli terms, the intermediate Pauli basis size is bounded by $2K$, ensuring that the overall computational scaling remains tractable throughout the simulation.
To assess the scalability of the algorithm under a fixed computational budget, we perform numerical experiments using the fixed-$K$ truncation scheme while increasing the system size $N$. We set $K$ equal to the number of Pauli terms generated for $N=12$ spins in the absence of truncation, namely K=2,704,156K=2{,}704{,}156. Using this fixed value of $K$, we simulate imaginary time evolution for system sizes ranging from $N=12$ to $N=50$ spins in steps of two. The Trotter step size is chosen as $\Delta t=0.04$, consistent with the previous experiments, and the evolution is carried out up to a maximum imaginary time $t=20.0$. The results are shown in Fig. 3.
From these simulations, we observe that a fixed-$K$ truncation yields reasonably accurate ground-state energies up to system sizes of approximately $N\simeq 20$. For larger systems, the ITPP energy saturates and no longer decreases appreciably with increasing $N$, indicating a breakdown of the approximation due to the imposed constraint on the Pauli basis size. Interestingly, this loss of accuracy coincides with the regime $K<2^{N}$, suggesting that the number of Pauli terms required to faithfully approximate the ground-state energy under imaginary time evolution may scale exponentially with system size. By contrast, in the intermediate regime $2^{N}<K<4^{N}$, ITPP maintains low relative error while retaining a number of Pauli terms that is substantially smaller than that of the full Pauli basis.
@@
As an additional consistency check, we performed simulations in which the truncation size was scaled exponentially with the system size for moderate system sizes, setting $K=2^{N}$ for $N$ ranging from $12$ to $24$ in steps of two. In this regime, both the absolute and relative errors remained approximately constant as the system size increased. Although these results are limited to a restricted range of system sizes, they provide further support for the interpretation that the loss of accuracy observed under fixed-$K$ truncation is primarily due to insufficient operator support during imaginary time evolution. More generally, these observations are consistent with the possibility that maintaining a comparable level of accuracy may require a number of Pauli terms that grows rapidly, potentially exponentially, with system size.
@@ IV Conclusions
In this work, we extend the Pauli Propagation framework to imaginary-time evolution. By deriving explicit update rules governing the propagation of Pauli operators under imaginary time operators generated by Pauli strings, we introduced an operator-based algorithm for approximating imaginary time evolution. This extension enables the direct computation of thermal and ground-state properties via imaginary time dynamics, while preserving the key computational advantages of Pauli Propagation, including the efficient bitstring representation of Pauli operators and the use of simple algebraic update rules.
@@
At a formal level, we demonstrated that Pauli Propagation can be consistently generalized to imaginary time evolution. In the absence of truncation, the resulting imaginary time Pauli Propagation algorithm (ITPP) exactly reproduces Trotterized imaginary time evolution, thereby providing an alternative operator-based formulation of imaginary time dynamics. We benchmarked ITPP on the one-dimensional transverse-field Ising model in the ordered phase, employing both coefficient-based and fixed-rank truncation strategies to control the growth of the Pauli basis.
@@
When truncation is introduced, ITPP yields approximate ground-state energies with controllable accuracy while retaining only a fraction of the full Pauli expansion. In particular, coefficient-threshold truncation enables a systematic trade-off between accuracy and computational cost, allowing meaningful ground-state energy estimates to be obtained with significantly fewer Pauli terms than required by the full, non-truncated evolution.
@@
At the same time, our numerical results highlight important limitations of Pauli-based imaginary time evolution. For coefficient-threshold truncation, the number of Pauli terms required to maintain a given level of accuracy grows rapidly with system size, while fixed number truncation experiments further show that, under a constant computational budget, ITPP accurately captures ground-state energies only up to moderate system sizes, beyond which the approximation deteriorates. Additional consistency checks, in which the truncation size was scaled exponentially with system size for small systems, suggest that maintaining comparable accuracy may require a exponentially growing number of Pauli terms. While these observations do not constitute a definitive characterization of the asymptotic scaling, they clearly illustrate the challenges associated with controlling operator growth in imaginary time Pauli Propagation while maintaining accurate ground-state energy approximations.
@@
Beyond closed-system dynamics, the extension to imaginary time evolution also suggests a natural pathway toward simulating open quantum system dynamics when combined with Pauli Propagation for real-time evolution. In principle, this can be achieved by working with the vectorized Lindblad equation, in which the evolution of the vectorized density matrix is generated by an operator containing both Hermitian (unitary) and anti-Hermitian (non-unitary) contributions. Within this framework, the unitary part can be treated using real-time Pauli Propagation, while the non-unitary part can be handled using the imaginary time extension introduced here. Related strategies have been explored for simulating open quantum dynamics on quantum hardware in Ref. [PRXQuantum.3.010320]. However, implementing this approach within the present framework would require a reformulation of Pauli Propagation in terms of vectorized operators rather than operator-valued observables, necessitating substantial modifications to the current software infrastructure. A vectorized Pauli-based approach has been investigated in Ref. [rudolph2023classicalsurrogatesimulationquantum], and a detailed exploration of this direction is left for future work.
@@
Taken together, our results establish ITPP as a useful and flexible tool for exploring imaginary time dynamics and ground-state properties in small to intermediate system sizes, while also clarifying the fundamental challenges posed by operator growth in non-unitary evolution. By making these limitations explicit, this work provides a clear baseline for future developments aimed at mitigating operator proliferation through improved truncation schemes or enhanced Pauli-based ansätze, as well as for extending Pauli Propagation to the simulation of open quantum system dynamics.
@@ V Acknowledgments
Appendix A Pauli Propagation Rules for imaginary time Evolution
In this appendix, we derive the propagation rules governing the imaginary time evolution of a Pauli string $P$ generated by another Pauli string $Q$. We consider the non-unitary imaginary time evolution operator
@@ Appendix A Pauli Propagation Rules for imaginary time Evolution
where $\tau\in\mathbb{R}$ denotes the imaginary time and $Q$ is a Pauli string satisfying $Q^{2}=\mathbb{I}$. The imaginary time–evolved Pauli operator is defined as
@@
The propagation of a Pauli string $P$ under imaginary time evolution generated by a Pauli string $Q$ is therefore given by
@@
These propagation rules form the basis of Pauli propagation under imaginary time evolution employed in imaginary time Pauli propagation (ITPP).

### v2

@@ III Numerical experiments
In this section, we present numerical results illustrating the performance of ITPP for approximating ground-state properties of the one-dimensional transverse-field Ising model (TFIM) with open boundary conditions. The Hamiltonian for $N$ spins is
$H_{\mathrm{TFIM}}=-J\sum_{i=1}^{N-1}Z_{i}Z_{i+1}-h\sum_{i=1}^{N}X_{i},$ (20)
@@
All simulations were performed using the PauliPropagation.jl Julia package [rudolph2025pauli], which provides an efficient implementation of Pauli propagation for real-time dynamics. We extend this framework by incorporating the imaginary-time propagation rules introduced in Sec. II.3. A related Julia package, PauliStrings.jl [Loizeau_2025], also provides tools for working with Pauli-string representations.
Our numerical study focuses on the ordered phase ($J>h$), using the parameter values $J=1.0$ and $h=0.5$. In this setting, we propagate the identity operator in imaginary time, according to Algorithm LABEL:alg:ITPP, and evaluate the resulting ground-state energy under various truncation schemes.
We first verify that, without truncation, ITPP faithfully reproduces Trotterized imaginary-time evolution and compare its performance with exact imaginary-time evolution obtained via exact diagonalization. Simulations are performed for a chain of $N=10$ spins with a Trotter step size $\Delta\tau=0.04$ and truncation threshold set to $\delta=0$. Throughout the imaginary-time evolution, we monitor the energy and its relative error with respect to the exact ground-state energy. The relative error is defined as
$\varepsilon_{E}(\tau)=\frac{\left|E(\tau)-E_{0}\right|}{\left|E_{0}\right|},$ (21)
where $E(\tau)$ is the energy at imaginary time $\tau$, and $E_{0}$ denotes the exact ground-state energy. The corresponding results are presented in Fig. 1.
From the results, we observe that ITPP exactly matches the Trotterized imaginary-time evolution, as expected, and closely follows the exact imaginary-time evolution until Trotter error becomes significant.
ITPP performance for $N=12$
Figure 2: Performance of ITPP with coefficient-threshold truncation for various thresholds in the ordered-phase TFIM ($J=1$, $h=0.5$) with $N=12$ spins and Trotter step $\Delta\tau=0.04$. The three panels show, from left to right, the energy during imaginary-time evolution, the relative error with respect to the exact ground-state energy (displayed on a logarithmic scale), and the number of Pauli terms retained under different truncation thresholds. In the left-most panel, the solid black line indicates the exact ground-state energy computed using the Bogoliubov–de Gennes solution. The results illustrate the trade-off between accuracy and computational cost controlled by the truncation threshold.
We next assess the performance of ITPP under a truncation strategy based on the magnitude of Pauli coefficients. The algorithm is executed using truncation thresholds ranging from $2^{-14}$ to $2^{-5}$, along with the idealized case of no truncation. Simulations are performed for a chain of $N=12$ spins with a Trotter step size $\Delta\tau=0.04$. Throughout the imaginary-time evolution, we track the energy, the relative error with respect to the exact ground-state energy, and the number of Pauli terms retained by the algorithm. The corresponding results are presented in Fig. 2.
@@
For thresholds larger than $2^{-7}$, however, we observe a drastic reduction in the number of Pauli terms retained during the evolution, which results in poor approximations of the true ground state. This behavior can be attributed to the tree-like structure underlying Pauli Propagation: during imaginary-time evolution, Pauli strings can generate many descendants at later times. If certain branches of this propagation tree are truncated too early, the algorithm discards Pauli terms that would otherwise contribute significantly at later stages. As a result, the evolution becomes confined to an overly restricted operator basis, leading to a loss of accuracy in the final ground-state approximation.
Truncation based on coefficient magnitude provides a natural way to discard Pauli terms by retaining only those with sufficiently large weights. However, this strategy does not offer direct control over the computational cost: fixing a threshold a priori does not determine how many Pauli terms will be retained during the evolution. In practice, the number of surviving terms can vary significantly depending on the system size and the stage of imaginary-time evolution.
ITPP with fixed-$K$ truncation and increasing NN
Figure 3: Scalability of ITPP with fixed-$K$ truncation for the one-dimensional TFIM in the ordered phase ($J=1$, $h=0.5$), for system sizes ranging from $N=12$ to $N=50$ spins in steps of two. The left panel shows the converged ground-state energy obtained with ITPP as a function of system size NN, while the right panel displays the corresponding relative error (on a logarithmic scale) with respect to the exact ground-state energy computed using the Bogoliubov–de Gennes solution. The truncation parameter is fixed to K=2,704,156K=2{,}704{,}156, equal to the number of Pauli terms generated for $N=12$ in the absence of truncation. The results indicate the range of system sizes for which this fixed Pauli basis size is sufficient to accurately capture the underlying operator growth under imaginary-time evolution, as well as the onset of a breakdown when the fixed computational budget becomes inadequate.
@@
We apply this fixed-$K$ truncation after each imaginary-time operator. Since each operation can generate at most twice as many Pauli terms, the intermediate Pauli basis size is bounded by $2K$, ensuring that the overall computational scaling remains tractable throughout the simulation.
To assess the scalability of the algorithm under a fixed computational budget, we perform numerical experiments using the fixed-$K$ truncation scheme while increasing the system size $N$. We set $K$ equal to the number of Pauli terms generated for $N=12$ spins in the absence of truncation, namely K=2,704,156K=2{,}704{,}156. Using this fixed value of $K$, we simulate imaginary-time evolution for system sizes ranging from $N=12$ to $N=50$ spins in steps of two. The Trotter step size is chosen as $\Delta\tau=0.04$, consistent with the previous experiments, and the evolution is carried out up to a maximum imaginary-time $\tau=20.0$. The results are shown in Fig. 3.
From these simulations, we observe that a fixed-$K$ truncation yields reasonably accurate ground-state energies up to system sizes of approximately $N\simeq 20$. For larger systems, the ITPP energy saturates and no longer decreases appreciably with increasing $N$, indicating a breakdown of the approximation due to the imposed constraint on the Pauli basis size. Interestingly, this loss of accuracy coincides with the regime $K<2^{N}$, suggesting that the number of Pauli terms required to faithfully approximate the ground-state energy under imaginary-time evolution may scale exponentially with system size. By contrast, in the intermediate regime $2^{N}<K<4^{N}$, ITPP maintains low relative error while retaining a number of Pauli terms that is substantially smaller than that of the full Pauli basis.
@@ IV Conclusions
In this work, we have extended the Pauli Propagation framework to imaginary-time evolution. By deriving explicit update rules governing the propagation of Pauli operators under imaginary-time operators generated by Pauli strings, we introduced an operator-based algorithm for approximating imaginary-time evolution. This extension enables the direct computation of thermal and ground-state properties, while preserving the key computational advantages of Pauli Propagation, including the efficient bitstring representation of Pauli operators and the use of simple algebraic update rules.
@@
At a formal level, we demonstrated that Pauli Propagation can be consistently generalized to imaginary-time evolution. In the absence of truncation, the resulting imaginary-time Pauli Propagation algorithm (ITPP) exactly reproduces Trotterized imaginary-time evolution, thereby providing an alternative operator-based formulation of imaginary-time dynamics. We benchmarked ITPP on the one-dimensional transverse-field Ising model in the ordered phase, employing both coefficient-based and fixed-rank truncation strategies to control the growth of the Pauli basis.
@@
When truncation is introduced, ITPP yields approximate ground-state energies with controllable accuracy while retaining only a small fraction of the full Pauli expansion. In particular, coefficient-threshold truncation enables a systematic trade-off between accuracy and computational cost, allowing meaningful ground-state energy estimates to be obtained with significantly fewer Pauli terms than required by the full, non-truncated evolution.
@@
At the same time, our numerical results highlight important limitations of Pauli-based imaginary-time evolution. For coefficient-threshold truncation, the number of Pauli terms required to maintain a given level of accuracy grows rapidly with system size, while fixed number truncation experiments further show that, under a constant computational budget, ITPP accurately captures ground-state energies only up to moderate system sizes, beyond which the approximation deteriorates. Additional consistency checks, in which the truncation size was scaled exponentially with system size for small systems, suggest that maintaining comparable accuracy may require an exponentially growing number of Pauli terms. While these observations do not constitute a definitive characterization of the asymptotic scaling, they clearly illustrate the challenges associated with controlling operator growth in ITPP while maintaining accurate ground-state energy approximations.
@@
The extension to imaginary-time evolution also suggests a natural pathway toward simulating more general open quantum system dynamics when combined with Pauli Propagation for real-time evolution. In principle, this can be achieved by working with the vectorized Lindblad equation, in which the evolution of the vectorized density matrix is generated by an operator containing both Hermitian (unitary) and anti-Hermitian (non-unitary) contributions. Within this framework, the unitary part can be treated using real-time Pauli Propagation, while the non-unitary part can be handled using the imaginary-time extension introduced here. Related strategies have been explored for simulating open quantum dynamics on quantum hardware in Ref. [PRXQuantum.3.010320]. However, implementing this approach within the present framework would require a reformulation of Pauli Propagation in terms of vectorized operators rather than operator-valued observables, necessitating substantial modifications to the current software infrastructure. A vectorized Pauli-based approach has been investigated in Ref. [rudolph2023classicalsurrogatesimulationquantum], and a detailed exploration of this direction is left for future work.
@@
Taken together, our results establish ITPP as a useful and flexible tool for exploring imaginary-time dynamics and ground-state properties in small to intermediate system sizes, while also clarifying the fundamental challenges posed by operator growth in non-unitary evolution. By making these limitations explicit, this work provides a clear baseline for future developments aimed at mitigating operator proliferation through improved truncation schemes or enhanced Pauli-based ansätze, as well as for extending Pauli Propagation to the simulation of open quantum system dynamics.
@@ V Acknowledgments
Appendix A Pauli Propagation Rules for Imaginary-Time Evolution
In this appendix, we derive the propagation rules governing the imaginary-time evolution of a Pauli string $P$ generated by another Pauli string $Q$. We consider the non-unitary imaginary-time evolution operator
@@ Appendix A Pauli Propagation Rules for Imaginary-Time Evolution
where $\tau\in\mathbb{R}$ denotes the imaginary-time and $Q$ is a Pauli string satisfying $Q^{2}=\mathbb{I}$. The imaginary-time–evolved Pauli operator is defined as
@@
The propagation of a Pauli string $P$ under imaginary-time evolution generated by a Pauli string $Q$ is therefore given by
@@
These propagation rules form the basis of Pauli propagation under imaginary-time evolution employed in ITPP.